# Exploiting Precision Mapping and Component-Specific Feature Enhancement for Breast Cancer Segmentation and Identification

### v4

@@ Abstract
Breast cancer is a leading cause of mortality worldwide, and demands the critical need for early and accurate diagnostic tools. Ultrasound imaging is a widely used modality for breast cancer screening, yet the precise segmentation and classification of tumors in these images are challenging due to variations in tumor morphology and image quality. To address these challenges, we propose novel deep learning (DL) frameworks leveraging a precision mapping mechanism (PMM) along with a component-specific feature enhancement module (CSFEM) to improve breast cancer lesion segmentation and identification. Our PPM ensures that the segmentation accurately reflects the true shape and extent of the tumor by meticulously delineating their boundaries. The CSFEM focuses on extracting and amplifying features unique to different tumor types, enabling the model to effectively distinguish between benign, malignant, and normal tissues. Integrating PMM and CSFEM into our segmentation model yielded an accuracy of 98.1%, an IoU of 96.9%, and a Dice Coefficient of 97.2%. Similarly, our classification model achieved an accuracy of 99.2%, with F1-score, precision, and recall values of 99.1%, 99.3%, and 99.1%, respectively. Our results indicate significant improvement in evaluation metrics in comparison to state-of-the-art (SOTA) models, demonstrating the effectiveness of precision mapping and component-specific feature enhancement in advancing breast cancer lesion analysis.
Keywords Breast Cancer  $\cdot$ Deep Learning  $\cdot$ Attention Mechanisms  $\cdot$ Medical Imaging
@@ 1 Introduction
Breast cancer is one of the most common cancers among women worldwide, resulting in approximately 570,000 deaths in 2015 alone. Annually, over 1.5 million women, accounting for 25% of all female cancer diagnoses, are diagnosed with breast cancer globally [1][2]. Breast tumors often originate as ductal hyperproliferation and can progress to benign tumors or metastatic carcinomas when stimulated by various carcinogenic agents. The tumor microenvironment, including stromal effects and macrophages, plays a crucial role in the development and progression of breast cancer [3].
@@
Early detection of breast carcinoma significantly increases the chances of successful treatment. Therefore, implementing effective procedures for identifying early signs of breast cancer is crucial [4]. Mammography, ultrasound, and thermography are the primary imaging techniques used for screening and diagnosing breast cancer [5][6]. With over 75% of tumors responding to hormones, breast cancer is primarily a postmenopausal illness. Their incidence rates are at the highest between the ages of 35-39 and then plateau after 80 years, with age and female sex being significant risk factors. This hormone dependency interacts with environmental and genetic factors to determine the incidence and progression of the disease [7].
Precise segmentation and classification of breast cancer are essential for effective treatment planning and positive patient outcomes. Traditional methods heavily depend on manual interpretation, which is both time-consuming and prone to errors. Advancements in technology have transformed the provision of healthcare. High processing power, primarily from GPUs, enables the creation of deep neural networks with multiple layers, allowing for the extraction of formerly unachievable features. Convolutional Neural Networks (CNNs) have made a profound impact on image processing and understanding, especially in the areas of segmentation, classification, and analysis [8][9].
@@
Deep learning models can process vast amounts of medical imaging data and detect subtle abnormalities that might elude human observers. Accurate tumor segmentation and classification enhances oncologists’ capacity to make decisions about whether a tumor is malignant or not. Typically, these methods require professional annotation and pathology reports to make this assessment [10], which consumes a lot of human effort. DL provides an efficient and promising solution for the automation of these procedures. They can learn complicated patterns and features from ultrasounds and mamograms, which has the potential to improve classification accuracy and efficiency.
This paper proposes the Spatial-Channel Attention LinkNet Framework with InceptionResNet Backbone for breast cancer segmentation, and DCNNIMAF Framework for breast cancer classification. The segmentation framework is a novel and effective attention-enhanced mechanism that uses a pre-trained CNN model architecture for the encoder backbone. This enhances the capability of feature extraction, while effectively enhancing segmentation using a coupled spatial and channel attention mechanism in the decoder. The proposed classification framework - Deep CNN with an Integrated Multi-Attention Framework (DCNNIMAF) - is a unique and novel architecture with a hybrid of integrated self and spatial attention mechanisms. The segmentation results were evaluated using evaluation metrics such as Dice coefficient, IoU score, and a combination of focal loss and Jaccard loss, while classification evaluation metrics include recall, F1-score, precision, and accuracy.
@@ 2 Related Works
Li et al. [12] introduced a novel patches screening method that included the extraction of multi-size and discriminative patches from histology images involving tissue-level and cell-level features. Firstly, patches of dimensions 512x512 and 128x128 are generated from the input data. This is followed by the utilization of two ResNet50s where one of the models is fed with patches of dimensions 128x128 while the other inputs patches of dimensions 512x512 which extract tissue-level and cell-level features respectively. A finetuning approach is adopted to train the ResNet50 models this is followed by a screening of patches by aggregating them into different clusters based on their phenotype. For speeding up the process, the patch size is reduced to obtain 1024 features followed by PCA to reduce the number of features to 200. This is followed by the k mean clustering process. A ResNet50 fine-tuned with 128x128 size patches is employed to select the clusters. Subsequently, the P-norm pooling feature method is applied to extract the final features of the image, followed by the use of a Support Vector Machine to classify input images into four distinct classes: Normal, Benign, In situ carcinoma, or Invasive carcinoma.
Zheng et al. [13] introduced a DL-assisted Efficient Adaboost Algorithm (DLA-EABA) where the Convolutional Neural Network is trained with extensive data so that high precision can be achieved. A stacked autoencoder is utilized for generating a deep convolutional neural network and the encoder and decoder sections contain multiple non-linear transformations which are taken from the combined depictions of actual data which is taken as input. An efficient Adaboost algorithm is utilized to train the classifiers which estimate the positive value for threshold and parity and is done by reviewing all the potential mixtures of both values, The deep CNN contains Long Short-Term Memory (LSTM) with logistic activation function as conventional artificial neurons. This is followed by Softmax Regression for classifying the images with the help of features extracted.
Lotter et al. [14] introduced a robust breast tumor classification model for mammography images which utilizes bounding box annotations and is extended to digital breast tomosynthesis images to be able to identify the tumor region in the image. The CNN first trains to classify if lesions are present in the cropped image patches. Subsequently, using the entire image as input, the CNN initializes the backbone of the detection-based model. This model outputs the entire image with a bounding box, providing a classification score. The model’s performance is then evaluated by comparing its ability to identify the tumor region with Breast Imaging Reporting and Data System Standard (BI-RADS) scores of 1 and 1 considered as negative interpretations and index and pre-index cancer exams.
@@
Saber et al. [15] employed transfer learning methodology on five different models: ResNet50, VGG19, Inception V3, Inception-V2, and VGG16. Feature extraction involved freezing the trained parameters from the source task except for the last three layers, which were then transferred to the target task. The images were preprocessed using different methods such as Median Filter, Histogram Equalization, Morphological Analysis, Segmentation, and Image Resizing. The dataset is split into an 80-20 ratio and Augmentation is applied to the training dataset where the images are rotated and flipped. The newly trained layers are combined with the existing pre-trained layers and the features are extracted using these models. Classification is done by feeding the extracted features from the transfer learning models into a Support Vector Machine classifier and Softmax classifiers that are fine-tuned using the Stochastic Gradient Descent method with momentum (SGDM). The gradient’s high-velocity dimensions are reduced due to SGDM jittering and the past gradients with momentum are reduced to saddle point.
@@
Cho et al. [16] proposed a Breast Tumor Ensemble Classification Network (BTEC-Net) which utilizes an improved DenseNet121 and ResNet101 as base classifiers where each of the four blocks is connected to the Squeeze and Excitation Block and Global Average Pooling layer. Next, the feature map sizes are aligned using a fully connected layer and integrated along the channel dimension. The combined feature map is then fed into a feature-level fusion module to perform binary classification. Once the classification is done, segmentation is carried out by utilizing the proposed Residual Feature Selection UNet model (RFS-UNet) which is an encoder-decoder network and are connected with the layer positions of the same feature map size using skip-connections. The encoder part is composed of five encoders with each one comprising of a convolutional layer, an RFS module, a residual convolutional block, and a max-pooling layer. Similarly, the model is composed of five decoders where each decoder comprises a convolutional layer and an RFS module as well, a transpose convolutional layer and a Residual Block. The skip connections contain a spatial attention module where the input involves the output of transposed convolution and output of the RFS module from the encoder and the output is concatenated to the output of the same transposed convolution layer. The segmentation process ends with a sigmoid activation function which returns the segmented tumor region.
Dayong Wang et al. [17] introduced a novel method for automatically detecting metastatic breast cancer in whole slide images of sentinel lymph node biopsies, achieving first place in the International Symposium on Biomedical Imaging (ISBI) grand challenge. Their system delivered impressive results with an AUC of 0.925 for whole slide image classification and a 0.7051 tumor localization score, surpassing an independent pathologist’s review. By integrating the DL system’s predictions with pathologist diagnoses, a notable reduction in the error rate was achieved, showcasing the profound impact of DL on enhancing the accuracy of pathological diagnoses for breast cancer metastases.
Abdelrahman Sayed Sayed et al. [18] developed a new, economical design for a 3-RRR Planar Parallel Manipulator (PPM), aiming to overcome the challenge of deriving kinematic constraint equations for manipulators with complex nonlinear behavior. Utilizing screw theory, they computed direct and inverse kinematics and then developed a Neuro-Fuzzy Inference System (NFIS) model that was optimized with Particle Swarm Optimization (PSO) and Genetic Algorithm (GA) to predict the position of the end-effector. The proposed PPM structure underwent investigation, with the development of its kinematic model and subsequent testing of a prototype in ADAMS, followed by fabrication for validation. Results showed that PSO outperformed GA in tuning the NFIS model, aligning closely with actual PPM data, indicating promise for enhanced robot capabilities and performance through further optimization and control strategies.
Luuk Balkenende et al. [19] proposed a comprehensive review elucidating the integration of deep learning techniques in breast cancer imaging. Their research highlights the wide-ranging applications of DL across modalities such as digital mammography, ultrasound, and magnetic resonance imaging (MRI), with a focus on tasks including lesion classification, segmentation, and predicting therapy response. Additionally, they discuss research on diagnosing breast cancer metastasis using CNNs on whole-body scintigraphy scans, and their investigation into aiding clinicians in diagnosing axillary lymph node metastasis with a 3D CNN model on PET/CT images. They emphasize the necessity of conducting large-scale trials and addressing ethical considerations to fully harness the potential of deep learning in clinical breast cancer imaging.
Shen et al. [20] proposed a pioneering DL-based approach for detecting breast cancer on screening mammograms. Their innovative "end-to-end" algorithm efficiently utilizes training datasets with varying levels of annotation, achieving exceptional performance compared to previous methods. On independent test sets from diverse mammography platforms, the proposed method achieves per-image AUCs ranging from 0.88 to 0.98, with sensitivities between 86.1% and 86.7%. Notably, the algorithm’s transferability across different mammography platforms is demonstrated, requiring minimal additional data for fine-tuning. These results emphasize the potential of deep learning to revolutionize breast cancer screening, offering more accurate and efficient diagnostic tools for clinical applications.
Han et al. [21] introduced a novel method for breast cancer diagnosis and prognosis. Their Class Structure-based Deep CNN (CSDCNN) achieves impressive accuracy (average 93.2%) by addressing challenges in automated multi-class classification from histopathological images. Combining hierarchical feature representation and distance constraints in feature space, their methodology offers a unique solution to subtle differences among breast cancer classes. Comparative experiments highlight the superior performance of the CSDCNN compared to existing methods, positioning it as a valuable tool for clinical decision-making in breast cancer management. Their work represents a significant advancement in automated breast cancer classification, providing clinicians with a reliable diagnostic aid.
Wang et al. [22] introduced DeepGrade, a deep learning-based histological grading model aimed at improving prognostic stratification for NHG 2 tumors. Developed and validated on large-scale datasets of digital whole-slide histopathology images, DeepGrade offers a novel approach to classify NHG 1 and NHG 3 morphological patterns. By re-stratifying NHG 2 tumors into DG2-high and DG2-low groups, DeepGrade provides independent prognostic information beyond traditional risk factors. Its performance was validated internally and externally, showcasing its ability to predict recurrence risk accurately. The ensemble approach, employing 20 deep convolutional neural network models, ensures robustness and reliability in classification tasks. DeepGrade shows promise as a cost-effective alternative to molecular profiling, supported by high area under the receiver operating characteristic curve values. This innovative methodology heralds a significant advancement in histological grading for breast cancer, promising improved clinical decision-making and personalized treatment strategies. Further research should focus on validating DeepGrade across diverse patient populations and integrating it into routine clinical practice.
Sizilio et al. [23] introduced a fuzzy logic-based approach for pre-diagnosing breast cancer from Fine Needle Aspirate (FNA) analysis. Addressing the global burden of breast cancer and the variability in FNA diagnostic accuracy (65% to 98%), this method enhances reliability through computational intelligence. The research employed the Wisconsin Diagnostic Breast Cancer Data (WDBC) and proceeded through four stages: fuzzification, rule base establishment, inference processing, and defuzzification. Validation included cross-validation and expert reviews. The method achieved a sensitivity of 98.59% and a specificity of 85.43%, demonstrating high reliability in detecting malignancies but highlighting the need for improvement in identifying benign cases. This approach shows significant potential for enhancing breast cancer diagnostic accuracy.
Sarkar et al. [24] explored the use of the K-Nearest Neighbors (KNN) algorithm for diagnosing breast cancer with the Wisconsin-Madison Breast Cancer dataset. Recognized for its straightforward and efficient implementation, KNN served as a non-parametric classifier in this study. The research showed that KNN improved classification performance by 1.17% over the best-known result for the dataset. Advantages of KNN include its simplicity, effectiveness with small training sets, and no need for retraining when new data is incorporated. However, the algorithm also has significant limitations, such as substantial storage requirements for large datasets and extensive computational demands for distance calculations between test and training data. The study noted the existence of faster KNN variants, such as those using k-d trees, which have been successful in tasks like script and speech recognition. The findings highlight KNN’s potential for various diagnostic applications, even though no single algorithm is optimal for all diagnostic problems. This research emphasizes KNN’s promise in enhancing diagnostic accuracy while acknowledging its challenges with storage and computational efficiency.
Song et al. [25] introduced an ML technique aimed at accurately annotating noncoding RNAs (ncRNAs) by searching genomes to find ncRNA genes characterized by known secondary structures. Their method involves aligning sequences optimally with a structure model, a critical step for identifying ncRNAs within genomes. Acknowledging the limitations of using a single structure model, they developed an approach that processes genome sequence segments to extract feature vectors. These vectors are then classified to differentiate between ncRNA family members and other sequences. The results showed that this method captures essential features of ncRNA families more effectively and enhances the accuracy of genome annotation compared to traditional tools. This work underscores the significant role of ML in bioinformatics, particularly in improving the precision of ncRNA gene identification.
Foster et al. [26] offered a critical commentary on the integration of ML in biomedical engineering, particularly focusing on the application of support vector machines (SVMs) beyond mere statistical tools. Their analysis highlighted the inherent challenges in developing clinically validated diagnostic techniques using SVMs, emphasizing concerns such as overfitting and the imperative for robust validation procedures. Unlike studies focused on specific diseases, their research aimed to evaluate and enhance existing ML models for broader biomedical applications. The commentary serves as a cautionary perspective for researchers, reviewers, and readers, stressing the complexities and potential pitfalls in classifier development. It advocates for an integrated approach where classifier validation forms an integral part of the experimental process. This work underscores the critical need to establish the clinical validity of diagnostic tools developed through ML in biomedical research.
Wei et al. [27] proposed an innovative method for improving microcalcification classification in breast cancer diagnosis using content-based image retrieval (CBIR) combined with ML. Their approach integrates CBIR to retrieve similar mammogram cases, enhancing the performance of a support vector machine (SVM) classifier. By incorporating local proximity information from retrieved cases, the adaptive SVM achieved a notable increase in classification accuracy from 78% to 82%, as measured by the area under the ROC curve. This method aims to provide radiologists with enhanced diagnostic support, serving as a valuable "second opinion" tool. Despite these advancements, the study acknowledges limitations in dataset size, which may affect generalizability. These findings underscore the potential of CBIR-assisted classification approaches in improving the precision of breast cancer diagnostics, emphasizing the need for further validation with larger clinical datasets to validate its efficacy and applicability in real-world clinical settings.

### v5

@@ Abstract
Breast cancer is one of the leading causes of death globally, and thus there is an urgent need for early and accurate diagnostic techniques. Although ultrasound imaging is a widely used technique for breast cancer screening, it faces challenges such as poor boundary delineation caused by variations in tumor morphology and reduced diagnostic accuracy due to inconsistent image quality. To address these challenges, we propose novel Deep Learning (DL) frameworks for breast lesion segmentation and classification. We introduce a precision mapping mechanism (PMM) for a precision mapping and attention-driven LinkNet (PMAD-LinkNet) segmentation framework that dynamically adapts spatial mappings through morphological variation analysis, enabling precise pixel-level refinement of tumor boundaries. Subsequently, we introduce a component-specific feature enhancement module (CSFEM) for a component-specific feature-enhanced classifier (CSFEC-Net). Through a multi-level attention approach, the CSFEM magnifies distinguishing features of benign, malignant, and normal tissues. The proposed frameworks are evaluated against existing literature and a diverse set of state-of-the-art Convolutional Neural Network (CNN) architectures. The obtained results show that our segmentation model achieves an accuracy of 98.1%, an IoU of 96.9%, and a Dice Coefficient of 97.2%. For the classification model, an accuracy of 99.2% is achieved with F1-score, precision, and recall values of 99.1%, 99.3%, and 99.1%, respectively.
Keywords Precision Mapping  $\cdot$ Component-Specific Feature Enhancement  $\cdot$ Multi-level Attention  $\cdot$ Medical Imaging
@@ 1 Introduction
Breast cancer is one of the most common cancers among women worldwide, causing about 570,000 deaths in 2015 alone. Every year, more than 1.5 million women, or 25% of all female cancer diagnoses, are diagnosed with breast cancer worldwide [1][2]. Breast tumors often begin as ductal hyperproliferation and can become benign tumors or metastatic carcinomas when stimulated by various carcinogenic agents. The tumor microenvironment, such as stromal effects and macrophages, contributes to the development and progression of breast cancer [3].
@@
Early detection of breast carcinoma increases the chances of successful treatment. Thus, there is a need for the implementation of effective procedures for the early detection of breast cancer signs [4]. Mammography, ultrasound, and thermography are some of the primary imaging techniques used in screening and diagnosing breast cancer [5][6]. With over 75% of the tumors being responsive to hormones, breast cancer is mostly a postmenopausal disease. Their incidence rates are at their peak ages between 35-39 years and then stabilize after the age of 80 years with age and female sex as major risk factors. This hormone dependence combines with environmental and genetic factors in determining the incidence and natural history of the disease [7]. Thus, accurate segmentation and classification of breast cancer are important for proper treatment planning and positive patient outcomes.
@@
The traditional methods rely a lot on manual interpretation that consumes a lot of time and has room for errors. Advancements in technology have transformed the provision of healthcare [8]. High processing power, primarily from GPUs, enables the creation of deep neural networks with multiple layers, allowing for the extraction of formerly unachievable features. CNNs have made a profound impact on image processing and understanding, especially in the areas of medical image segmentation, classification, and analysis [9][10]. DL models can process vast amounts of medical imaging data and detect subtle abnormalities that might elude human observers. Accurate tumor segmentation and classification enhances oncologists’ capacity to make decisions about whether a tumor is malignant or not. Typically, these methods require professional annotation and pathology reports to make this assessment [11], which consumes a lot of human effort. DL provides an efficient and promising solution for the automation of these procedures [12]. They can learn complicated patterns and features from ultrasounds and mamograms, which has the potential to improve classification accuracy and efficiency.
We introduce, in this paper, a novel Precision Mapping and Attention-Driven LinkNet framework, with an integrated Precision Mapping Mechanism (PMM) module. PMM dynamically calibrates pixel-level mappings to achieve accurate boundary delineation, compensating for morphological variations and inconsistencies in medical images. A specially designed attention-driven architecture, leveraging a pre-trained encoder backbone, optimizes feature extraction. The decoder uses a coupled spatial-channel attention mechanism to dynamically refine spatial features with enhanced semantic context for accurate tumor segmentation. We also propose a Component-Specific Feature Enhancement Module (CSFEM) to develop a Component-Specific Feature-Enhanced classifier - CSFEC-Net. Selective amplification of discriminative patterns allows the CSFEM to dynamically learn and amplify tissue-specific characteristics that improve feature representation, and enhance inter-class separability. Segmentation results were validated by employing the Dice coefficient, IoU score, and a combination of focal loss with Jaccard loss. On the other hand, classification is evaluated using the following evaluation metrics: recall, F1-score, precision, and accuracy.
Our proposed contributions are:
  • Precision Mapping Mechanism (PMM): A novel boundary refinement module for a PMAD-LinkNet segmentation framework that integrates an adaptive mapping strategy with pixel-level precision to localize tumor edges with high accuracy. This is done by dynamically calibrating spatial mappings according to morphological variations and compensating for inconsistencies in image quality through a learned refinement process.
  • Component-Specific Feature Enhancement Module (CSFEM): A dedicated feature augmentation module for CSFEC-Net that uses dynamic multi-attention mechanisms to selectively amplify discriminative traits of benign, malignant, and normal tissues. Using a multi-scale attention pipeline, CSFEM enhances inter-class separability and intra-class consistency and thus can effectively address the problem of overlapping feature spaces in heterogeneous medical imagery.
@@ 2 Related Works
Li et al. [14] introduced a novel patches screening method that included the extraction of multi-size and discriminative patches from histology images involving tissue-level and cell-level features. Firstly, patches of dimensions 512x512 and 128x128 are generated from the input data. This is followed by the utilization of two ResNet50s where one of the models is fed with patches of dimensions 128x128 while the other inputs patches of dimensions 512x512 which extract tissue-level and cell-level features respectively. A finetuning approach is adopted to train the ResNet50 models this is followed by a screening of patches by aggregating them into different clusters based on their phenotype. For speeding up the process, the patch size is reduced to obtain 1024 features followed by PCA to reduce the number of features to 200. This is followed by the k mean clustering process. A ResNet50 fine-tuned with 128x128 size patches is employed to select the clusters. Subsequently, the P-norm pooling feature method is applied to extract the final features of the image, followed by the use of a Support Vector Machine to classify input images into four distinct classes: Normal, Benign, In situ carcinoma, or Invasive carcinoma. Zheng et al. [15] introduced a DL-assisted Efficient Adaboost Algorithm (DLA-EABA) where the Convolutional Neural Network is trained with extensive data so that high precision can be achieved. A stacked autoencoder is utilized for generating a deep convolutional neural network and the encoder and decoder sections contain multiple non-linear transformations which are taken from the combined depictions of actual data which is taken as input. An efficient Adaboost algorithm is utilized to train the classifiers which estimate the positive value for threshold and parity and is done by reviewing all the potential mixtures of both values, The deep CNN contains Long Short-Term Memory (LSTM) with logistic activation function as conventional artificial neurons. This is followed by Softmax Regression for classifying the images with the help of features extracted.
@@
Lotter et al. [16] introduced a robust breast tumor classification model for mammography images which utilizes bounding box annotations and is extended to digital breast tomosynthesis images to be able to identify the tumor region in the image. The CNN first trains to classify if lesions are present in the cropped image patches. Subsequently, using the entire image as input, the CNN initializes the backbone of the detection-based model. This model outputs the entire image with a bounding box, providing a classification score. The model’s performance is then evaluated by comparing its ability to identify the tumor region with Breast Imaging Reporting and Data System Standard (BI-RADS) scores of 1 and 1 considered as negative interpretations and index and pre-index cancer exams. Saber et al. [17] employed transfer learning methodology on five different models: ResNet50, VGG19, Inception V3, Inception-V2, and VGG16. Feature extraction involved freezing the trained parameters from the source task except for the last three layers, which were then transferred to the target task. The images were preprocessed using different methods such as Median Filter, Histogram Equalization, Morphological Analysis, Segmentation, and Image Resizing. The dataset is split into an 80-20 ratio and Augmentation is applied to the training dataset where the images are rotated and flipped. The newly trained layers are combined with the existing pre-trained layers and the features are extracted using these models. Classification is done by feeding the extracted features from the transfer learning models into a Support Vector Machine classifier and Softmax classifiers that are fine-tuned using the Stochastic Gradient Descent method with momentum (SGDM). The gradient’s high-velocity dimensions are reduced due to SGDM jittering and the past gradients with momentum are reduced to saddle point.
@@
Cho et al. [18] proposed a Breast Tumor Ensemble Classification Network (BTEC-Net) which utilizes an improved DenseNet121 and ResNet101 as base classifiers where each of the four blocks is connected to the Squeeze and Excitation Block and Global Average Pooling layer. Next, the feature map sizes are aligned using a fully connected layer and integrated along the channel dimension. The combined feature map is then fed into a feature-level fusion module to perform binary classification. Once the classification is done, segmentation is carried out by utilizing the proposed Residual Feature Selection UNet model (RFS-UNet) which is an encoder-decoder network and are connected with the layer positions of the same feature map size using skip-connections. In the encoder part, there are five encoders and each one includes a convolutional layer, an RFS module, a residual convolutional block and a max-pooling layer. The model also consists of five decoders in which each decoder contains a convolutional layer and an RFS module, a transpose convolutional layer and a Residual Block as well. The skip connections are equipped with a spatial attention module which takes as input the output of the transposed convolution and the output of the RFS from the encoder, and produces concatenation of outputs to the output of the same transposed convolution layer. The last layer in the segmentation process uses a sigmoid activation function which retrieves the output which is the segmented tumor region. Dayong Wang et al.[19] presented a new approach for detecting metastatic breast cancer in whole-slide images of sentinel lymph node biopsies and won the first place in the ISBI 2016 grand challenge. Their system performed satisfactorily with an AUC of 0.925 for whole-slide image classification and 0.7051 as a tumor localization score that surpassed an independent pathologist’s review. Combining predictions of the DL system and diagnostics of pathologists, a significant decrease of the error rate was obtained.
Abdelrahman Sayed Sayed et al. [20] proposed a novel low-cost structure of a 3-RRR Planar Parallel Manipulator (PPM) in their research – attempting to address the problem of formulating kinematic constraint equations for drive manipulators with complex nonlinear dependence. They performed direct and inverse kinematics techniques using screw theory and established a Neuro-Fuzzy Inference System (NFIS) model optimized with Particle Swarm Optimization (PSO) and Genetic Algorithm (GA) for end-effector position prediction. Proposed PPM structure was emphasized, including kinematic modeling and preliminary PPM virtual tests within ADAMS, followed by verification prototype manufacture. Results demonstrated PSO as a more successful tuning approach for the NFIS model than GA, which provided better consistency with the experimental PPM data and thus the strategy holds potential for further robot enhancement and performance improvement opportunities from additional optimization and control measures. A thorough review by Luuk Balkenende et al. [21] described an integrated application of deep learning techniques in imaging breast cancer. Their paper demonstrates a variety of applications in digital mammography, ultrasound colonography, and magnetic resonance imaging and introduces some applications such as lesion classification, segmentation, and therapy response prediction. The research is also highlighted by findings on the autonomous diagnosis of breast cancer metastasis with convolutional neural networks on whole-body scintigraphy scans and the investigation into helping clinicians diagnose axillary lymph node metastases using a 3D CNN model on PET/CT images. They mention the necessity to conduct large-scale trials and address ethical considerations in order to realize the full potential of deep learning in clinical breast cancer imaging.
A novel approach to the detection of breast cancer in screening mammograms based on deep learning was presented by Shen et al. [22]. This "end-to-end" algorithm is claimed to use efficiently training sets of various levels of annotation to perform really well by comparison with previous methods. Across the independent test sets collected from different mammography platforms, the method reaches per-image AUCs ranging from 0.88 to 0.98, coupled with sensitivities in the range of 86.1%-86.7%. Most important is that it has demonstrated cross-platform transferability, which requires only a minimal amount of additional data for fine-tuning. The results underlined the promise of deep learning in transforming breast cancer screening into more accurate and efficient diagnostic tools and thus into clinical applications. A novel approach for breast cancer diagnosis and prognosis employing a class structure-based Deep CNN was established by Han et al [23]. In automated multi-class classification from histopathological images, their CSDCNN addresses the caveat and does an impressive average accuracy of 93.2%. By establishing a hierarchical feature representation with distance constraints in feature space, this methodology serves as a unique avenue to minor differences between the various classes of breast cancer to which comparative experiment results are significantly favorable for the CSDCNN compared to other methods. It also poises itself as one of the most valuable clinical decision aids in breast cancer management. This study has provided an important impetus in the direction of automated breast cancer classification and reliable diagnoses aid to clinicians.
Sizilio et al. [24] proposed a fuzzy logic approach to pre-diagnosing breast cancer from FNA analysis. The method was introduced in the context of the global burden of breast cancer and also a wide range of accuracy in FNA-based diagnosis, from 65 to 98 percent. Hence, this method increases the reliability of computing intelligence. The study made use of the Wisconsin Diagnostic Breast Cancer Data (WDBC) and followed fuzzification, rule-base creation, inference processing, and defuzzification. Validation was based on cross-validation and expert reviews. The method achieved a sensitivity of 98.95% and specificity of 85.43%, implying a very reliable malignant detection but a very poor benign detection. This technique has a good potential for improving diagnostic accuracy in breast cancer. The application of K-Nearest Neighbours (KNN) for breast cancer diagnosis with the Wisconsin-Madison Breast Cancer dataset was explored by Sarkar et al. [25]. KNN was relatively straightforward and efficient in its non-parametric classifier use in this respect. This research showed that KNN increased the classification performance with an improvement of 1.17% over the best-known result for the dataset. KNN is simple, effective for smaller training sets, and it does not require retraining of trained models when new data is available. However, limitations also arise as it incurs huge storage overhead for larger datasets and high computation requirements to calculate distances between test and training data. The study mentions that there are faster KNN variants, such as k-d trees, which have performed excellently in script and speech recognition tasks. KNN has much more potential in regard to diagnostic applications, even though one algorithm is not suitable for all problems in diagnostics. This research states not only the promise of KNN in improving the accuracy of diagnostics but also addresses the challenges of storage and computational efficiency.
Foster et al. [26] provided a critical overview of the role of ML in biomedical engineering, paying attention to the extension of SVMs beyond the use of this statistical tool. Their study revealed the intrinsic challenges in constructing SVM-based clinically validated diagnostic systems, pointing out problems, such as overfitting and the need for thorough validation protocols. Unlike disease-specific studies, their study was designed to evaluate and enhance existing ML models for a broader biomedical application. The commentary is a cautionary view for researchers, reviewers, and readers in emphasizing the complexities and pitfalls that may accompany classifier development. It supports an integrated approach where validation of classifiers is part of the experimental process. This work puts great emphasis on the fact that, in biomedical research, clinical validity should be ascertained for the diagnosis tools developed by ML. Wei et al. [27] introduced an innovative approach which improves the classification performance of the proposed diagnostic system based on the classification of breast cancer images by microcalcifications, which is developed through a combination of CBIR and ML. The combination of CBIR improves the retrieval of similar mammogram cases for enhanced SVM classifier performance. Utilizing recovered case information with local proximity, the adaptive SVM achieved significantly better classification accuracy compared with the baseline model at 78% to 82%, as measured in the area under the ROC curve. The technique could make for a good source of support for radiologists while delivering an enhanced "second opinion" tool. Still, the study underlines problems with small datasets that may well influence generalization ability. These results suggest that assistive classification approaches using CBIR may improve the accuracy in the diagnosis of breast cancers, but further validation, possibly with larger clinical data sets, will be required to confirm such efficacy and practical applicability in actual clinical settings.